# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ III-H Optimizing The RBS MAs’ Position
In this subsection, with other variables being given, we investigate the update of the posotion of the $m$-th RBS’ $n$-th MA, i.e., $\mathbf{t}^{1}_{m,n}$. Firstly, the function $\tilde{\mathrm{R}}_{u,l}$ given in (30) is rewritten as
$\!-\!{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}\!=\!{\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{7}_{m,l,n,n}\mathbf{h}^{1}_{m,l,n}$ (72)
$\!+\!\!\textrm{Re}\{(\mathbf{d}^{9}_{m,l,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}\!\!+\!\!\!{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{11}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}\!\!+\!\!c_{5,1,m,n},$
with $c_{5,1,m,n}$ being a constant and the parameters in (72) are defined as follows in (73).
$\mathbf{D}^{1}_{m,n,l,i}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H})^{T}\otimes(q_{i}(\mathbf{\Sigma}^{1}_{n,i})^{H}\mathbf{h}^{0,1}_{n,i}\mathbf{h}^{0,1}_{m,i}\mathbf{\Sigma}^{1}_{m,i}),\hat{\mathbf{h}}^{1}_{m,l}\triangleq\textrm{vec}(\mathbf{H}^{1}_{m,l}(\mathbf{t}^{1}_{m})),\mathbf{d}^{1}_{m,l,i}\triangleq 2({\sum}_{j\neq m}^{M_{t}}(\mathbf{D}^{1}_{m,j,l,i})^{H}\hat{\mathbf{h}}^{1}_{j,i}),$ (73)
$\mathbf{d}^{2}_{m,l}\triangleq 2\omega_{u,l}\beta_{u,l}\sqrt{q_{l}}\mathbf{u}_{m,l}^{H}\otimes((\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{\Sigma}^{1}_{m,l}),\mathbf{D}^{2}_{j,l}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H})],\mathbf{D}^{4}_{j,l}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{D}^{2}_{j,l}+\mathbf{D}^{3}_{j,l}),$
$\mathbf{D}^{3}_{j,l}\triangleq(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}),\bar{\mathbf{D}}^{4}_{j,l,m,n}\triangleq\mathbf{D}^{4}_{j,l}[(m-1)N_{r}+1:mN_{r},(n-1)N_{r}+1:nN_{r}],$
$\mathbf{d}^{3}_{j,l,m}\triangleq 2({\sum}_{i\neq m}^{M_{r}}(\bar{\mathbf{D}}^{4}_{j,l,i,m})^{H}\mathbf{g}_{r,i,j}),\mathbf{D}^{5}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{D}^{1}_{m,m,l,i},\mathbf{d}^{4}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{1}_{m,l,i}-\mathbf{d}^{2}_{m,l},\mathbf{D}^{6}_{j,m}\triangleq{\sum}_{l=1}^{K_{u}}\bar{\mathbf{D}}^{4}_{j,l,m,m},$
$\mathbf{d}^{6}_{j,m}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{d}^{3}_{j,l,m},\mathbf{D}^{7}_{m,l,n,j}\triangleq\mathbf{D}^{5}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(j-1)L^{t,1}_{m,l}+1:jL^{t,1}_{m,l}],\mathbf{d}^{8}_{m,l,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}(\mathbf{D}^{7}_{m,l,n,j})^{H}\mathbf{h}^{1}_{m,l,n}),$
$\mathbf{d}^{9}_{m,l,n}\triangleq\mathbf{d}^{8}_{m,l,n}+\mathbf{d}^{4}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l}],d^{10}_{j,m,n}\triangleq 2({\sum}_{i\neq n}^{N_{r}}(\mathbf{D}^{6}_{j,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])+\mathbf{d}^{6}_{j,m}[n].$
Furthermore, the sensing SINR constraint (28b) can be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\Longleftrightarrow{\sum}_{l=1}^{K_{u}}\big{(}(\mathbf{h}^{1}_{m,l,n})^{H}\mathbf{D}^{15}_{l,m,n,n}\mathbf{h}^{1}_{m,l,n}$ (74)
$\!+\!\!\textrm{Re}\{(\mathbf{d}^{16}_{l,m,n})^{H}\mathbf{h}^{1}_{m,l,n}\}\big{)}+{\sum}_{j=1}^{K_{t}}\big{(}\textrm{Re}\{(d^{17}_{j,m,n})^{\ast}\mathbf{g}_{r,m,j}[n]\}\big{)}$
$-\textrm{Re}\{(d^{19}_{m,n})^{\ast}\mathbf{g}_{r,m,0}[n]\}+c_{5,2,m,n}\leq 0,$
@@
where $c_{5,2,m,n}$ is a constant term and the above newly introduced coefficients given in (75).
$\mathbf{D}^{10}\triangleq\big{(}{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]+(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})\big{)}/\Gamma_{t},$ (75)
$\bar{\mathbf{D}}^{10}_{n,m}\triangleq\mathbf{D}^{10}[(n-1)N_{r}+1:nN_{r},(m-1)N_{r}+1:mN_{r}],\mathbf{D}^{11}_{l,m,n}\triangleq(\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H})^{T}\otimes(q_{l}(\mathbf{\Sigma}^{1}_{n,l})^{H}\mathbf{h}^{0,1}_{n,l}\mathbf{h}^{0,1}_{m,l}\Sigma^{1}_{m,l}),$
$\mathbf{d}^{12}_{l,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}(\mathbf{D}^{11}_{l,m,n})^{H}\hat{\mathbf{h}}^{1}_{n,l}),\mathbf{D}^{14}_{j}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]+(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}),$
$\bar{\mathbf{D}}^{14}_{j,i,z}\triangleq\mathbf{D}^{14}_{j}[(i-1)N_{r}+1:iN_{r},(z-1)N_{r}+1:zN_{r}],\mathbf{d}^{13}_{j,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,j}^{H}\bar{\mathbf{D}}^{14}_{j,n,m})),\mathbf{d}^{14}_{m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,0}^{H}\bar{\mathbf{D}}^{10}_{n,m})),$
${\mathbf{D}}^{15}_{l,m,n,i}\triangleq\mathbf{D}^{11}_{l,m,m}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(i-1)L^{t,1}_{m,l}+1:iL^{t,1}_{m,l}],\mathbf{d}^{15}_{l,m,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}({\mathbf{D}}^{15}_{l,m,j,n})^{H}\mathbf{h}^{1}_{m,l,j}),$
$\mathbf{d}^{16}_{l,m,n}\!\triangleq\!\mathbf{d}^{15}_{l,m,n}\!+\!\mathbf{d}^{12}_{l,m}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],d^{17}_{j,m,n}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{14}_{j,m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{13}_{j,m}[n],$
$d^{19}_{m,n}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{10}_{m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,0}[i])\!+\!\mathbf{d}^{14}_{m}[n].$

### v3

@@ III-H Optimizing The RBS MAs’ Position
In this subsection, with other variables being given, we investigate the update of the position of the $m$-th RBS’ $n$-th MA, i.e., $\mathbf{t}^{1}_{m,n}$  [^7] To improve readability, we have removed the subscripts, i.e., $m$ and $n$, from the variables and/or coefficients. Specifically, we represent: $\mathbf{t}^{1}_{m,n}$ as $\mathbf{t}^{1}$, $\mathbf{h}^{1}_{m,l,n}$ as $\mathbf{h}^{1}_{l}$, and $\mathbf{g}_{r,m,j}[n]$ as ${g}_{r,j}$. .
Firstly, the function $\tilde{\mathrm{R}}_{u,l}$ given in (32) is rewritten as
$-{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}={\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{1}_{l}\mathbf{h}^{1}_{l}+\textrm{Re}\{(\mathbf{d}^{1}_{l})^{H}\mathbf{h}^{1}_{l}\}$ (77)
$-\textrm{Re}\{(\mathbf{{d}}^{2}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)+{\sum}_{j=0}^{K_{t}}\textrm{Re}\{(d^{3}_{j})^{\ast}\mathbf{g}_{r,j}\}+c_{5,1},$
with $c_{5,1}$ being a constant and the parameters in (77) defining in following
$\mathbf{\bar{D}}^{1}_{m,n,l,i}\!\!\triangleq\!\omega_{u,l}|\beta_{u,l}|^{2}\!(\mathbf{u}_{m,l}\mathbf{u}_{n,l}^{H})^{T}\!\!\!\otimes\!\!(q_{i}(\mathbf{\Sigma}^{1}_{n,i})^{H}\!\mathbf{h}^{0,1}_{n,i}\!\mathbf{h}^{0,1}_{m,i}\mathbf{\Sigma}^{1}_{m,i}),$ (78)
$\hat{\mathbf{h}}^{1}_{m,l}\triangleq\textrm{vec}(\mathbf{H}^{1}_{m,l}(\mathbf{t}^{1}_{m})),\mathbf{d}^{19}_{m,l,i}\triangleq 2({\sum}_{j\neq m}^{M_{t}}(\mathbf{\bar{D}}^{1}_{m,j,l,i})^{H}\hat{\mathbf{h}}^{1}_{j,i}),$
$\mathbf{d}^{20}_{m,l}\triangleq 2\omega_{u,l}\beta_{u,l}\sqrt{q_{l}}\mathbf{u}_{m,l}^{H}\otimes((\mathbf{h}^{0,1}_{m,l})^{H}\mathbf{\Sigma}^{1}_{m,l}),$
$\mathbf{\bar{D}}^{2}_{j,l}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H})],$
$\mathbf{D}^{4}_{j,l}\triangleq\omega_{u,l}|\beta_{u,l}|^{2}(\mathbf{\bar{D}}^{2}_{j,l}+\mathbf{D}^{3}_{j,l}),$
$\mathbf{D}^{3}_{j,l}\triangleq(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{l}\hat{\mathbf{u}}_{l}^{H}),$
$\bar{\mathbf{D}}^{4}_{j,l,m,n}\triangleq\mathbf{D}^{4}_{j,l}[(m-1)N_{r}+1:mN_{r},(n-1)N_{r}+1:nN_{r}],$
$\mathbf{d}^{21}_{j,l,m}\triangleq 2({\sum}_{i\neq m}^{M_{r}}(\bar{\mathbf{D}}^{4}_{j,l,i,m})^{H}\mathbf{g}_{r,i,j}),\mathbf{D}^{5}_{m,i}\triangleq{\sum}_{l=1}^{K_{u}}\mathbf{\bar{D}}^{1}_{m,m,l,i},$
$\mathbf{d}^{22}_{m,i}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\mathbf{d}^{19}_{m,l,i},\mathbf{D}^{6}_{j,m}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\bar{\mathbf{D}}^{4}_{j,l,m,m},\mathbf{d}^{23}_{j,m}\!\triangleq\!\!{\sum}_{l=1}^{K_{u}}\mathbf{d}^{21}_{j,l,m},$
$\mathbf{D}^{1}_{l}\triangleq\mathbf{D}^{5}_{m,l}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(j-1)L^{t,1}_{m,l}+1:jL^{t,1}_{m,l}],$
$\mathbf{d}^{24}_{m,l,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}(\mathbf{D}^{7}_{m,l,n,j})^{H}\mathbf{h}^{1}_{m,l,n}),$
$\mathbf{d}^{1}_{l}\!\triangleq\!\mathbf{d}^{24}_{m,l,n}\!\!+\!\mathbf{d}^{22}_{m,l}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],$
$d^{3}_{j}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\mathbf{D}^{6}_{j,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{23}_{j,m}[n],$
$\mathbf{{d}}^{2}_{l}\!\!\triangleq\!\mathbf{d}^{20}_{m,l}[(n\!-\!1)L^{t,1}_{m,l}+1\!:\!nL^{t,1}_{m,l}].$
And the sensing SINR constraint (30b) can be rewritten as
$\textrm{SINR}_{t}\geq\Gamma_{r}\!\!\!\Longleftrightarrow\!\!\!{\sum}_{l=1}^{K_{u}}\big((\mathbf{h}^{1}_{l})^{H}\mathbf{D}^{2}_{l}\mathbf{h}^{1}_{l}\!+\!\!\textrm{Re}\{(\mathbf{d}^{4}_{l})^{H}\mathbf{h}^{1}_{l}\}\big)$ (79)
$+{\sum}_{j=1}^{K_{t}}\big(\textrm{Re}\{(d^{5}_{j})^{\ast}\mathbf{g}_{r,j}\}\big)-\textrm{Re}\{(d^{6})^{\ast}{g}_{r,0}\}+c_{5,2}\leq 0,$
@@
where $c_{5,2}$ is a constant term and the above newly introduced coefficients are given in
$\mathbf{D}^{10}\triangleq\big({\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]$ (80)
$+(\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,0})^{T}\otimes\sigma^{2}_{t,0}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})\big)/\Gamma_{t},$
$\bar{\mathbf{D}}^{10}_{n,m}\triangleq\mathbf{D}^{10}[(n-1)N_{r}+1:nN_{r},(m-1)N_{r}+1:mN_{r}],$
$\mathbf{D}^{11}_{l,m,n}\triangleq(\mathbf{u}_{m,0}\mathbf{u}_{n,0}^{H})^{T}\otimes(q_{l}(\mathbf{\Sigma}^{1}_{n,l})^{H}\mathbf{h}^{0,1}_{n,l}\mathbf{h}^{0,1}_{m,l}\Sigma^{1}_{m,l}),$
$\mathbf{d}^{25}_{l,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}(\mathbf{D}^{11}_{l,m,n})^{H}\hat{\mathbf{h}}^{1}_{n,l}),$
$\mathbf{D}^{14}_{j}\triangleq{\sum}_{k=1}^{K_{d}}[(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}\hat{\mathbf{w}}_{k}^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H})]$
$+(\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}^{r}(\hat{\mathbf{W}}^{r})^{H}\hat{\mathbf{g}}_{t,j})^{T}\otimes\sigma^{2}_{t,j}(\hat{\mathbf{u}}_{0}\hat{\mathbf{u}}_{0}^{H}),$
$\bar{\mathbf{D}}^{14}_{j,i,z}\triangleq\mathbf{D}^{14}_{j}[(i-1)N_{r}+1:iN_{r},(z-1)N_{r}+1:zN_{r}],$
$\mathbf{d}^{26}_{j,m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,j}^{H}\bar{\mathbf{D}}^{14}_{j,n,m})),\mathbf{d}^{27}_{m}\triangleq 2({\sum}_{n\neq m}^{M_{r}}({\mathbf{g}}_{r,n,0}^{H}\bar{\mathbf{D}}^{10}_{n,m})),$
${\mathbf{D}}^{2}_{l}\triangleq\mathbf{D}^{11}_{l,m,m}[(n-1)L^{t,1}_{m,l}+1:nL^{t,1}_{m,l},(i-1)L^{t,1}_{m,l}+1:iL^{t,1}_{m,l}],$
$\mathbf{d}^{28}_{l,m,n}\triangleq 2({\sum}_{j\neq n}^{N_{r}}({\mathbf{D}}^{15}_{l,m,j,n})^{H}\mathbf{h}^{1}_{m,l,j}),$
$\mathbf{d}^{4}_{l}\!\triangleq\!\mathbf{d}^{28}_{l,m,n}\!+\!\mathbf{d}^{25}_{l,m}[(n\!-\!1)L^{t,1}_{m,l}\!+\!1\!:\!nL^{t,1}_{m,l}],$
$d^{5}_{j}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{14}_{j,m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,j}[i])\!+\!\mathbf{d}^{26}_{j,m}[n],$
$d^{6}\!\triangleq\!2({\sum}_{i\neq n}^{N_{r}}(\bar{\mathbf{D}}^{10}_{m,m}[i,n])^{\ast}{\mathbf{g}}_{r,m,0}[i])\!+\!\mathbf{d}^{27}_{m}[n].$